From the knowledge that CV, the molar specific heat at constant volume, for a gas in a container is 5.0R, calculate the ratio of the speed of sound in that gas to the rms speed of the molecules,


for gas temperature T.

1
Expert's answer
2022-04-22T13:53:40-0400

"\frac v{v_s}=\sqrt{\frac\gamma 3}=1.3."